Write five pairs of prime numbers less than 20 whose sum is divisible by 5. (Hint: 3 + 7 = 10)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The five pairs of prime numbers less than 20 whose sum is divisible by 5 are
2 +3 =5
2 + 13 = 15
3 + 17 = 20
7 + 13 = 20
19 + 11 = 30

Thanks